OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. L(5, 0), M(0, 5)

OpenStudy (ivettef365):

First find the slope

OpenStudy (jack1):

m = y2 -y1 / x2 - x1 M = (x1,y1) L = (x2, y2)

OpenStudy (anonymous):

idk how...

OpenStudy (uri):

@jperez18 You should know the formula of slope first m = y2 -y1 / x2 - x1

OpenStudy (anonymous):

uri, that gave me -1. is that my answer?

OpenStudy (ivettef365):

now you do the equation y = mx + b where m = slope and b = y-int

OpenStudy (uri):

Do you know that :